OpenStudy (anonymous):

{2x+5y=-4 {4x-5y=22 what is the solution to the above system of equation

OpenStudy (anonymous):

look at the y's.. if you add these two equations together the y's will cancel out. 2x+5y=-4 4x-5y=22 + ------------ = 6x+0y=18 now x=18/6 = 3. now sub 3 back into any of the two equations to find y, so 2(3)+5y=-4 6+5y=-4 5y=-10 y=-10/5 = -2

OpenStudy (anonymous):

thank you so much!!!!

OpenStudy (anonymous):

this is both, you can start with eliminating one variable